# Magnetic-field induced momentum-dependent symmetry breaking in CsV3Sb5 revealed by magneto-ARPES

### v1

@@ Abstract
In quantum materials with multiple degrees of freedom with similar energy scales, intertwined electronic orders with distinct broken symmetries often appear in a strongly coupled fashion. Recently, in a class of kagome superconductors represented by CsV3Sb5, experimental reports have suggested rotational symmetry breaking and time reversal symmetry breaking associated with a charge density wave (CDW) order, revealing an exotic nature of this CDW order. Here, utilizing our recently developed capability of performing angle-resolved photoemission spectroscopy in a tunable magnetic field (magneto-ARPES), we reveal momentum-selective response of the electronic structure of CsV3Sb5 to an external magnetic field. While the response in the electronic structure is clearly compatible with piezomagnetism, strong orbital-selectivity is observed. Specifically, bands associated with the vanadium -orbitals contributing to the Van Hove singularities (VHS) near the Brillouin zone (BZ) boundary exhibit selective spectral broadening that breaks C6 rotational symmetry and is odd in magnetic field, disappearing above the CDW transition. Meanwhile, the antimony -orbital dominated electron pocket at the BZ center becomes elongated under an applied field—an effect that persists above the . Our observations delineate the origin of the time-reversal symmetry breaking associated with the vanadium VHS-bands at the onset of the CDW order, while the field-induced rotational symmetry breaking largely associated with the antimony orbitals reflects fluctuations beyond the CDW ordering temperature. Our magneto-ARPES work demonstrates a novel tuning knob for disentangling intertwined orders in the momentum space for quantum materials.
@@ I Main
The kagome lattice, a prototypical two-dimensional geometrically frustrated system, has been extensively studied as a promising candidate for realizing the quantum spin liquid state Syozi1951; Balents2010a; Han2012. Beyond spin interactions, the kagome lattice also exhibits rich physics associated with the quantum interference of the electronic hopping paths, hosting flat bands, Dirac cones, and Van Hove singularities (VHS) in its electronic structure Mielke1991; Tanaka2003; Yu2012a; Kiesel2012; Kiesel2013. When the flat bands or the VHSs are positioned near the Fermi level, electronic instabilities emerge even for relatively weak electron-electron interactions, potentially leading to the emergence of different symmetry-broken phases, such as magnetism, charge density wave (CDW), superconductivity, and electronic nematicity Norman2016; Ko2009; Yu2012a; Kiesel2012; Kiesel2013; Wang2013a. Such a proximity to competing ordered states makes the kagome systems highly susceptible to external perturbations such as strain and magnetic field, which can be used to tune the competition among the exotic phases Liu2020a; Wang2023a; Lima2023; Dey2011. Consequently, kagome metals provide an excellent platform for investigating the various intriguing symmetry-breaking phases due to the intricate interweaving among electron correlations, topology, and geometric frustration Balents2010a; Ko2009; Kang2020a.
@@
CsV3Sb5 has been identified as a kagome superconductor (Tc 3.2 K) with the V atoms forming the kagome lattice Ortiz2019; Ortiz2020. It undergoes a 222 CDW transition at 94 K, which is well captured by first-principles calculations, indicating the relevant role of the electron-phonon coupling Tan2021; Christensen2021. Interestingly, the CDW phase was also found to display unconventional properties suggestive of time-reversal symmetry-breaking (TRSB), which has been interpreted in terms of an interaction-driven loop-current phase Park2021; Lin2021; Denner2021; Christensen2022; Fischer2023; Tazai2024. These properties include an enhanced relaxation rate measured by muon spin rotation Mielke2022b; Khasanov2022, magnetic-field switching of the relative intensity of the CDW Bragg peaks seen in scanning tunneling microscopy Jiang2021; Xing2024, and anisotropic and non-reciprocal electrical magneto-transport Guo2022; Wei2024; Guo2024a. On the other hand, the presence of a non-zero magnetization remains under debate, in view of contradictory reports on the existence of a spontaneous Kerr effect Xu2022; Farhang2023; Saykin2023; Wang2024. Additionally, sixfold (C6) to twofold (C2) rotational symmetry breaking (RSB) was detected at around 35 K in the CDW state of CsV3Sb5 by a combination of nuclear magnetic resonance and elastoresistance measurements Nie2022a. Such a RSB was attributed to the interlayer stacking of the 222 CDW order Christensen2021 and interpreted as an electronic nematic orderXiang2021; Jin2024. While the interpretation of RSB in terms of nematic order has been challenged by recent studies that did not find a divergent nematic susceptibility Liu2024, various experimental probes in CsV3Sb5, including polarization-resolved Raman spectroscopy Wulferding2022, scanning tunneling microscopy Zhao2021a; Chen2021a; Li2022a, and optical spectroscopy Xu2022, have consistently found RSB within the CDW state. Interestingly, resistivity anisotropy Guo2024a and magnetic torque measurements Asaba2024 in CsV3Sb5 have proposed a connection between RSB and TRSB. Thus, to advance the understanding of this material and shed light on these conflicting experimental reports, a direct momentum-resolved spectroscopy measurement in the presence of a magnetic field is desirable.
@@
In light of this, we utilize our recently developed magneto-ARPES technique Huang2023a to directly measure the response of the momentum-resolved electronic structure of CsV3Sb5 to an out-of-plane tunable magnetic field. ARPES has been remarkably successful in directly probing the momentum-resolved electronic spectra of quantum materials, providing a way to disentangle intertwined order parameters that exhibit distinct momentum-dependent fingerprints Damascelli2003; Sobota2021. For the AV3Sb5 (A = K, Rb, Cs) compounds, ARPES studies have revealed the key electronic characteristics inherent to the kagome lattice, along with the band folding and CDW gap associated with the CDW transition Liu2021; Luo2022; Kang2022; Hu2022a. However, to date, no momentum-resolved electronic response to an external magnetic field has been reported using ARPES.
@@
Magnetic field is commonly used as a tuning knob for other types of experiments, but it has been carefully avoided in ARPES measurements due to its unpredictable effects on the photoelectron trajectory after photoemission Ryu2023a; Huang2023a. Recently, we have developed a simple method for implementing an in situ tunable out-of-plane magnetic field as the sample environment during ARPES measurements via a solenoid coil (Fig. 1e and f), enabling magneto-ARPES–a term coined by an independent work developing an in-plane field for ARPES Ryu2023a. While the magnetic fields available in our setup are small, below 10 mT, they are larger than the internal magnetic field values reported by SR in AV3Sb5 Mielke2022b; Khasanov2022. Utilizing this capability, we performed magneto-ARPES measurements on CsV3Sb5 to probe a momentum-dependent electronic response to a magnetic field. Specifically, we observed spectral modifications to the dispersions and Fermi surfaces associated with the vanadium -orbital VHS bands around the K/K’ points of the BZ. Importantly, these spectral modifications break the C6 symmetry of the dispersion and are odd in the magnetic field, consistent with piezomagnetism Xing2024. Additionally, we observe an elliptical elongation of the antimony -orbital electron pocket around the point of the BZ in the presence of a magnetic field. In contrast to the field-induced anisotropic spectral changes observed in the vanadium -bands, which disappear above , the spectral changes around the point persists above , suggesting significant fluctuations. Our magneto-ARPES results unambiguously reveal time-reversal symmetry breaking in CsV3Sb5 interwoven with rotational symmetry breaking, and a qualitatively different magnetic-field response of the low-energy electronic states arising from the Sb -orbitals and the V -orbitals. These results not only provide strong constraints to theoretical delineation of the exotic CDW order parameter in CsV3Sb5, but also establish a new capability for disentangling momentum-dependent order parameters in the presence of a field.
@@ II Field response of the V 3 VHS bands at K/K’
We first introduce the zero-field electronic structure of CsV3Sb5 measured using a helium lamp-based ARPES system (Fig. 1b and c). Consistent with previous reports, the Fermi surface of CsV3Sb5 is composed of a circular electron pocket around the BZ center consisting mainly of the 5 orbitals of Sb and three triangles around the BZ corner (K/K’) consisting of the 3 orbitals of V and related to the nearby VHS at the M points Kang2022; Hu2022a. The measured band dispersions match well with those obtained from density-functional theory (DFT) calculations (Fig. 1c). Raw magneto-ARPES data generally consist of intrinsic effects occurring inside the sample in response to the magnetic field and extrinsic effects occurring when the field acts on the photoemitted electrons on their way to the analyzer after they are ejected from the sample. The extrinsic effects consist of a constant energy contour rotation due to the Lorentz force, photoelectron emission angle contraction, and momentum broadening (Fig. 1d-f) Huang2023a. Specifically, for the small magnetic field adopted in this work (1.6 mT), only a constant energy contour rotation was discernible while the other two effects were negligible (Fig. 1g and h) Huang2023a. This in-plane constant energy contour rotation is effectively energy independent within the energy range examined in this work (up to 1 eV below the Fermi level, see Supplementary Materials), allowing straightforward post-measurement correction without the need for intricate data processing. It can also be used to directly determine the field strength at the sample. As an example, Fig. 1g-i compares the Fermi surface mapped by rotating the sample measured at 0, -1.6 mT and +1.6 mT. The maps display opposite rotation directions for positive and negative fields, which can be corrected by an in-plane azimuthal angle offset before any further data analysis is performed (Fig. 1h,i). To avoid any potential complications in field-dependent comparisons arising from the photoemission matrix-element effects Damascelli2003, all subsequent measurements were performed with the DA30 deflector mode without the need to rotate or move samples when changing the magnetic field. Hence, in all comparisons that we provide with and without a field, the beam spot was kept on a fixed spot of the sample during the whole measurement, guaranteeing that the light polarization relative to the crystal axes remained the same when the field strength was changed.
@@
To better illustrate this field-induced anisotropic spectral intensity change, spectral images along cuts across both K and K’ are extracted for a side-by-side comparison (Fig. 2e-h and Extended Data Fig. 1-2). The cuts are taken perpendicular to the –M mirror plane, where a mirror-symmetric spectral weight distribution is expected when there is no magnetic field. We note that the experimental geometry is chosen to preserve this mirror, as neither our beam polarization nor incident direction breaks this mirror symmetry (Fig. 2). We focus on the band as its spectral weight is discernible over a large energy range. For cut 1 within the first BZ, the ARPES spectra of the band on the left side () and those of the band on the right side () of the = 0 mirror plane are equivalent in the absence of a magnetic field (Fig. 2e). However, when an out-of-plane magnetic field (1.6 mT) is applied, the left branch of the band intensity becomes "weak" and the spectral weight becomes broader, as observed from the momentum distribution curve (MDC) in Fig. 2f. The spectral broadening is reversed with respect to the horizontal mirror plane when the magnetic field direction is reversed (Fig. 2g). This can be better seen in the magneto-dichroic ARPES signal shown in Fig. 2h, which we define as the difference in spectral intensity between the two spectra obtained in magnetic fields with opposite directions and the same strength. This magneto-dichroic signal can be well simulated by a simple selective momentum broadening of ARPES spectra based on the DFT calculated band structures along this cut (Extended Data Fig. 3). For cut 2, which lies outside the first BZ on the other side of the K–M–K BZ boundary, an opposite selective spectral weight broadening behavior is observed between and in a magnetic field compared with what is observed along cut 1 (Extended Data Fig. 1). The spectral evolution of the images along both cut 1 and cut 2 in an external magnetic field together complements the spectral response of the "X"-shaped energy iso-surface to a magnetic field, which directly breaks the –M mirror symmetry of the electronic states. We emphasize that this observation is an intrinsic electronic response of CsV3Sb5 to the external magnetic field, as the extrinsic momentum broadening effect is negligible at this small field and should be neither momentum selective nor dependent on the sign of the field Huang2023a.
@@
To pinpoint the precise symmetries that are broken by this anisotropic effect, we note that our measurement region also covers part of the second BZ, which can be mapped back to the first BZ via the crystalline translational symmetry. We also directly verified this by rotating the sample and mapping the different K/K’ pairs response to the magnetic field on the same sample (Extended Data Fig. 4). This leads to the illustration in Fig. 2i, where the broadened branches of the Fermi surface are depicted by the thicker and lighter-colored lines while the shaper branches are represented by the thinner and darker-colored lines. Comparing the behavior of the branches across the first BZ, we conclude that the inversion symmetry is preserved, whereas the sixfold (C6) rotational symmetry is broken. Moreover, since the effect changes sign when the field is reversed, the RSB is odd in field, which is indicative of piezomagnetism, i.e. a shear strain that is linear with the field Xing2024.
@@
The odd-in-field response of the electronic structure to the out-of-plane field suggests signs of TRS-breaking. To understand the origin of the observed asymmetric spectral shape under a magnetic field, we carried out one-step model ARPES calculations using the spin-polarized relativistic Korringa-Kohn-Rostoker (SPRKKR) package to examine the effect of an external magnetic field on the photoemission spectral function on CsV3Sb5 Ebert2011Calculating (see Methods and Extended Data Fig. 5). By calculating the spectral function of the non-CDW phase of CsV3Sb5 under an external field, we can simulate the photoemission spectra under a finite field for the simplest phase with broken TRS via the Zeeman effect without invoking a specific microscopic TRS-broken order. This set of calculations incorporates the magnetic field effects in all three parts of the photoemission matrix element—a Zeeman term for the initial state, the relativistic dipole operator, and the time-reversed spin-polarized low-energy electron diffraction final states, and is hence the simplest yet most unbiased way to examine the expectations from a generic TRS-broken state. As shown in Extended Data Fig. 5, the spectra near the K/K’ points exhibit asymmetric spectral shape both in peak height and widths when photoemitting under a large magnetic field. This is due to the combination of the field-induced band shift and spin-dependent multiple scattering in the photoemission process under a particular field direction. While at first sight this may seem consistent with our observations, the magnitude of the field needed for the Zeeman effect to be observable is five orders of magnitude larger than the field used in the magneto-ARPES measurements. Hence the electronic response experimentally observed here cannot arise from a Zeeman-like field-induced effect on the photoemission matrix elements, but must originate from an intrinsic TRS-breaking order parameter in CsV3Sb5.

### v2

# Magnetic field-induced momentum-dependent symmetry breaking in a kagome superconductor

@@ Abstract
When multiple degrees of freedom share similar energy scales in quantum materials, intertwined electronic orders, which exhibit broken symmetries, are often strongly coupled. Recent studies on kagome superconductors such as CsV3Sb5 report rotational and time-reversal symmetry breaking linked to a charge density wave. Here, we observe a momentum-selective response of the electronic structure of CsV3Sb5 to an external magnetic field. By performing angle-resolved photoemission spectroscopy in a tuneable magnetic field, we demonstrate that the response of the electronic structure is compatible with piezomagnetism along with strong orbital selectivity. Our results show that the origin of the time-reversal symmetry breaking is associated with the vanadium Van Hove singularities at the onset of the charge density wave order. We also demonstrate the presence of fluctuations beyond the charge ordering temperature. Our results reveal that magnetic fields can be used as tuning knobs for disentangling intertwined orders in the momentum space for quantum materials.
@@ I Main
The kagome lattice, a prototypical two-dimensional geometrically frustrated system, has been extensively studied as a promising candidate for realizing the quantum spin liquid state Syozi (1951); Balents (2010); Han et al. (2012). Beyond spin interactions, the kagome lattice also exhibits rich physics associated with the quantum interference of the electronic hopping paths, hosting flat bands, Dirac cones, and Van Hove singularities (VHS) in its electronic structure Mielke (1991); Tanaka and Ueda (2003); Yu and Li (2012); Kiesel and Thomale (2012); Kiesel et al. (2013). When the flat bands or the VHSs are positioned near the Fermi level, electronic instabilities emerge even for relatively weak electron-electron interactions, potentially leading to the emergence of different symmetry-broken phases, such as magnetism, charge density wave (CDW), superconductivity, and electronic nematicity Norman (2016); Ko et al. (2009); Yu and Li (2012); Kiesel and Thomale (2012); Kiesel et al. (2013); Wang et al. (2013). Such a proximity to competing ordered states makes the kagome systems highly susceptible to external perturbations such as strain and magnetic field, which can be used to tune the competition among the exotic phases Liu (2020); Wang et al. (2023); Lima et al. (2023); Dey et al. (2011). Consequently, kagome metals provide an excellent platform for investigating the various intriguing symmetry-breaking phases due to the intricate interweaving among electron correlations, topology, and geometric frustration Balents (2010); Ko et al. (2009); Kang et al. (2020).
@@
CsV3Sb5 has been identified as a kagome superconductor (Tc 3.2 K) with the V atoms forming the kagome lattice Ortiz et al. (2019, 2020). It undergoes a 222 CDW transition at 94 K, which is well captured by first-principles calculations, indicating the relevant role of the electron-phonon coupling Tan et al. (2021); Christensen et al. (2021). Interestingly, the CDW phase was also found to display unconventional properties suggestive of time-reversal symmetry-breaking (TRSB), which has been interpreted in terms of an interaction-driven loop-current phase Park et al. (2021); Lin and Nandkishore (2021); Denner et al. (2021); Christensen et al. (2022); Wagner et al. (2023); Tazai et al. (2024). These properties include an enhanced relaxation rate measured by muon spin rotation Mielke et al. (2022); Khasanov et al. (2022), magnetic-field switching of the relative intensity of the CDW Bragg peaks seen in scanning tunneling microscopy Jiang et al. (2021); Xing et al. (2024), and anisotropic and non-reciprocal electrical magneto-transport Guo et al. (2022); Wei et al. (2024); Guo et al. (2024b). On the other hand, the presence of a non-zero magnetization remains under debate, in view of contradictory reports on the existence of a spontaneous Kerr effect Xu et al. (2022); Farhang et al. (2023); Saykin et al. (2023); Wang et al. (2024). Additionally, sixfold (C6) to twofold (C2) rotational symmetry breaking (RSB) was detected at around 35 K in the CDW state of CsV3Sb5 by a combination of nuclear magnetic resonance and elastoresistance measurements Nie et al. (2022). Such a RSB was attributed to the interlayer stacking of the 222 CDW order Christensen et al. (2021) and interpreted as an electronic nematic orderXiang et al. (2021); Jin et al. (2024). While the interpretation of RSB in terms of nematic order has been challenged by recent studies that did not find a divergent nematic susceptibility Liu et al. (2024), various experimental probes in CsV3Sb5, including polarization-resolved Raman spectroscopy Wulferding et al. (2022), scanning tunneling microscopy Zhao et al. (2021); Chen et al. (2021); Li et al. (2022), and optical spectroscopy Xu et al. (2022), have consistently found RSB within the CDW state. Interestingly, resistivity anisotropy Guo et al. (2024b) and magnetic torque measurements Asaba et al. (2024) in CsV3Sb5 have proposed a connection between RSB and TRSB. Thus, to advance the understanding of this material and shed light on these conflicting experimental reports, a direct momentum-resolved spectroscopy measurement in the presence of a magnetic field is desirable.
@@
In light of this, we utilize our recently developed magneto-ARPES technique Huang et al. (2023) to directly measure the response of the momentum-resolved electronic structure of CsV3Sb5 to an out-of-plane tunable magnetic field. ARPES has been remarkably successful in directly probing the momentum-resolved electronic spectra of quantum materials, providing a way to disentangle intertwined order parameters that exhibit distinct momentum-dependent fingerprints Damascelli et al. (2003); Sobota et al. (2021). For the AV3Sb5 (A = K, Rb, Cs) compounds, ARPES studies have revealed the key electronic characteristics inherent to the kagome lattice, along with the band folding and CDW gap associated with the CDW transition Liu et al. (2021); Luo et al. (2022); Kang et al. (2022); Hu et al. (2022). However, to date, no momentum-resolved electronic response to an external magnetic field has been reported using ARPES.
@@
Magnetic field is commonly used as a tuning knob for other types of experiments, but it has been carefully avoided in ARPES measurements due to its unpredictable effects on the photoelectron trajectory after photoemission Ryu et al. (2023); Huang et al. (2023). Recently, we have developed a simple method for implementing an in situ tunable out-of-plane magnetic field as the sample environment during ARPES measurements via a solenoid coil (Fig. 1e and f), enabling magneto-ARPES–a term coined by an independent work developing an in-plane field for ARPES Ryu et al. (2023). While the magnetic fields available in our setup are small, below 10 mT, they are larger than the internal magnetic field values reported by SR in AV3Sb5 Mielke et al. (2022); Khasanov et al. (2022). Utilizing this capability, we performed magneto-ARPES measurements on CsV3Sb5 to probe a momentum-dependent electronic response to a magnetic field. Specifically, we observed spectral modifications to the dispersions and Fermi surfaces associated with the vanadium -orbital VHS bands around the K/K’ points of the BZ. Importantly, these spectral modifications break the C6 symmetry of the dispersion and are odd in the magnetic field, consistent with piezomagnetism Xing et al. (2024). Additionally, we observe an elliptical elongation of the antimony -orbital electron pocket around the point of the BZ in the presence of a magnetic field. In contrast to the field-induced anisotropic spectral changes observed in the vanadium -bands, which disappear above , the spectral changes around the point persists above , suggesting significant fluctuations. Our magneto-ARPES results unambiguously reveal time-reversal symmetry breaking in CsV3Sb5 interwoven with rotational symmetry breaking, and a qualitatively different magnetic-field response of the low-energy electronic states arising from the Sb -orbitals and the V -orbitals. These results not only provide strong constraints to theoretical delineation of the exotic CDW order parameter in CsV3Sb5, but also establish a new capability for disentangling momentum-dependent order parameters in the presence of a field.
@@ II Field response of the V 3 VHS bands at K/K’
We first introduce the zero-field electronic structure of CsV3Sb5 measured using a helium lamp-based ARPES system (Fig. 1b and c). Consistent with previous reports, the Fermi surface of CsV3Sb5 is composed of a circular electron pocket around the BZ center consisting mainly of the 5 orbitals of Sb and three triangles around the BZ corner (K/K’) consisting of the 3 orbitals of V and related to the nearby VHS at the M points Kang et al. (2022); Hu et al. (2022). The measured band dispersions match well with those obtained from density-functional theory (DFT) calculations (Fig. 1c). Raw magneto-ARPES data generally consist of intrinsic effects occurring inside the sample in response to the magnetic field and extrinsic effects occurring when the field acts on the photoemitted electrons on their way to the analyzer after they are ejected from the sample. The extrinsic effects consist of a constant energy contour rotation due to the Lorentz force, photoelectron emission angle contraction, and momentum broadening (Fig. 1d-f) Huang et al. (2023). Specifically, for the small magnetic field adopted in this work (1.6 mT), only a constant energy contour rotation was discernible while the other two effects were negligible (Fig. 1g and h) Huang et al. (2023). This in-plane constant energy contour rotation is effectively energy independent within the energy range examined in this work (up to 1 eV below the Fermi level, see Supplementary Materials), allowing straightforward post-measurement correction without the need for intricate data processing. It can also be used to directly determine the field strength at the sample. As an example, Fig. 1g-i compares the Fermi surface mapped by rotating the sample measured at 0, -1.6 mT and +1.6 mT. The maps display opposite rotation directions for positive and negative fields, which can be corrected by an in-plane azimuthal angle offset before any further data analysis is performed (Fig. 1h,i). To avoid any potential complications in field-dependent comparisons arising from the photoemission matrix-element effects Damascelli et al. (2003), all subsequent measurements were performed with the DA30 deflector mode without the need to rotate or move samples when changing the magnetic field. Hence, in all comparisons that we provide with and without a field, the beam spot was kept on a fixed spot of the sample during the whole measurement, guaranteeing that the light polarization relative to the crystal axes remained the same when the field strength was changed.
@@
To better illustrate this field-induced anisotropic spectral intensity change, spectral images along cuts across both K and K’ are extracted for a side-by-side comparison (Fig. 2e-h and Extended Data Fig. 1-2). The cuts are taken perpendicular to the –M mirror plane, where a mirror-symmetric spectral weight distribution is expected when there is no magnetic field. We note that the experimental geometry is chosen to preserve this mirror, as neither our beam polarization nor incident direction breaks this mirror symmetry (Fig. 2). We focus on the band as its spectral weight is discernible over a large energy range. For cut 1 within the first BZ, the ARPES spectra of the band on the left side () and those of the band on the right side () of the = 0 mirror plane are equivalent in the absence of a magnetic field (Fig. 2e). However, when an out-of-plane magnetic field (1.6 mT) is applied, the left branch of the band intensity becomes "weak" and the spectral weight becomes broader, as observed from the momentum distribution curve (MDC) in Fig. 2f. The spectral broadening is reversed with respect to the horizontal mirror plane when the magnetic field direction is reversed (Fig. 2g). This can be better seen in the magneto-dichroic ARPES signal shown in Fig. 2h, which we define as the difference in spectral intensity between the two spectra obtained in magnetic fields with opposite directions and the same strength. This magneto-dichroic signal can be well simulated by a simple selective momentum broadening of ARPES spectra based on the DFT calculated band structures along this cut (Extended Data Fig. 3). For cut 2, which lies outside the first BZ on the other side of the K–M–K BZ boundary, an opposite selective spectral weight broadening behavior is observed between and in a magnetic field compared with what is observed along cut 1 (Extended Data Fig. 1). The spectral evolution of the images along both cut 1 and cut 2 in an external magnetic field together complements the spectral response of the "X"-shaped energy iso-surface to a magnetic field, which directly breaks the –M mirror symmetry of the electronic states. We emphasize that this observation is an intrinsic electronic response of CsV3Sb5 to the external magnetic field, as the extrinsic momentum broadening effect is negligible at this small field and should be neither momentum selective nor dependent on the sign of the field Huang et al. (2023).
@@
To pinpoint the precise symmetries that are broken by this anisotropic effect, we note that our measurement region also covers part of the second BZ, which can be mapped back to the first BZ via the crystalline translational symmetry. We also directly verified this by rotating the sample and mapping the different K/K’ pairs response to the magnetic field on the same sample (Extended Data Fig. 4). This leads to the illustration in Fig. 2i, where the broadened branches of the Fermi surface are depicted by the thicker and lighter-colored lines while the shaper branches are represented by the thinner and darker-colored lines. Comparing the behavior of the branches across the first BZ, we conclude that the inversion symmetry is preserved, whereas the sixfold (C6) rotational symmetry is broken. Moreover, since the effect changes sign when the field is reversed, the RSB is odd in field, which is indicative of piezomagnetism, i.e. a shear strain that is linear with the field Xing et al. (2024).
@@
The odd-in-field response of the electronic structure to the out-of-plane field suggests signs of TRS-breaking. To understand the origin of the observed asymmetric spectral shape under a magnetic field, we carried out one-step model ARPES calculations using the spin-polarized relativistic Korringa-Kohn-Rostoker (SPRKKR) package to examine the effect of an external magnetic field on the photoemission spectral function on CsV3Sb5 Ebert et al. (2011) (see Methods and Extended Data Fig. 5). By calculating the spectral function of the non-CDW phase of CsV3Sb5 under an external field, we can simulate the photoemission spectra under a finite field for the simplest phase with broken TRS via the Zeeman effect without invoking a specific microscopic TRS-broken order. This set of calculations incorporates the magnetic field effects in all three parts of the photoemission matrix element—a Zeeman term for the initial state, the relativistic dipole operator, and the time-reversed spin-polarized low-energy electron diffraction final states, and is hence the simplest yet most unbiased way to examine the expectations from a generic TRS-broken state. As shown in Extended Data Fig. 5, the spectra near the K/K’ points exhibit asymmetric spectral shape both in peak height and widths when photoemitting under a large magnetic field. This is due to the combination of the field-induced band shift and spin-dependent multiple scattering in the photoemission process under a particular field direction. While at first sight this may seem consistent with our observations, the magnitude of the field needed for the Zeeman effect to be observable is five orders of magnitude larger than the field used in the magneto-ARPES measurements. Hence the electronic response experimentally observed here cannot arise from a Zeeman-like field-induced effect on the photoemission matrix elements, but must originate from an intrinsic TRS-breaking order parameter in CsV3Sb5.